# Unlearning regularization for Boltzmann Machines

### v2

@@ V The Hebbian Unlearning limit
We will now study the specific limit of the unlearning regularization that leads to an algorithm which strongly resembles the HU routine. The inferential performance of the learning procedure is examined by measuring the distance between the inferred model and the original one at different training steps. Results show, for the first time, that HU can be employed as an inferential tool, due to the beneficial effect of the Hebbian initialization of the couplings. We advance an explanation for its performance that is supported by further numerical evidence. In the limit $a\rightarrow 0$ the updating rules for the parameters transform into Eqs.  (28) and (29). As mentioned before, Eq. (28) strongly resembles the traditional HU algorithm of Ref. [8]. By contrast with the original rule, the new regularization samples configurations at $\beta=1$ instead of stable fixed points of the neural dynamics, and makes use of a thermal average, rather than summing each contribution each time. We will keep dealing with small networks so that, given the initial conditions for the parameters, the loss function of the problem can be minimized exactly. As a numerical experiment, we train a BM with $N=18$ neurons to learn a SK model at $\beta_{d}=0.4$. The unlearning regularization is applied with $a=0$ and the usual initial conditions for the parameters, i.e. $\bm{J}^{(0)}=\bm{c}_{d}$ and $h_{i}^{(0)}=0$ $\forall i$.
@@
VI Conclusions
@@
The goal of this article was thus to show the effectiveness of unlearning as a form of regularization in Boltzmann Machines. For this purpose, we have used small networks that allowed to compute all quantities without error by exact enumeration. Precisely, the ground-truth distribution was defined by one of two models, either the Curie-Weiss or the Sherrington-Kirkpatrick models, which are known to show a phase transition at a critical temperature. As a result, we could not only test the proximity of the inferred model to the original one, but also the susceptibility of the network under variations of the parameters, by comparing the trend of the specific heat of the ground truth model with respect to the inverse temperature, with the trend obtained after inference with regularization. We conclude that, given a high consistency between the model and the ground-truth distribution, the stability of the model under rescaling of the parameters is enhanced by the unlearning regularization. As an alternative strategy to the unlearning regularization presented here, we also evaluated the minimization of the KL distance between the models at $\beta=1$ and $\beta=a$. Even if the gradient descent equations look similar to the current case, they contain the contribution of four-point correlations among spins, resulting in a significant increase of the computational cost of the training.
We believe that the unlearning procedure can be employed in more realistic scenarios, where the ground-truth distribution is not known a-priori, through performing a cross-validation procedure. The real data-set should be then divided in two pieces: a training set, that is used to train the model with a certain choice of the control parameter $a$; a testing set, with an empirical data distribution, that is now compared to the inferred model at temperature $\beta=1$. The training procedure is repeated for several choices of $a$. One can choose to stop training earlier as soon as the slope of $\bm{c}_{d}$ vs. $\bm{c}_{1}$ approaches unity, or to let it run until convergence. The value of the variable $a$ that gives the highest resemblance of the model with the statistics of the testing set, as well as the lowest $C_{v}(1)$, is adopted as a winning choice for the hyper-parameter. Performing a test on a real data-set is a clear direction for future work.
@@
Moreover we analyzed the particular limit of the regularization that gives as learning rules the ones expressed in equations (28) and (29). The resulting procedure strongly resembles the traditional Hebbian Unlearning routine, but with a thermal two-point correlation computed at $\beta=1$ replacing that computed on the zero-temperature fixed points of the dynamics. Previous work [12] supports the fact that the associative memory performance of such a thermal unlearning does not deviate much from the original rule. Unlearning is not new to be mentioned in the statistical inference realm [34, 35, 33], yet we are the first to consider the original rule proposed by Hopfield in [8]. As a new contribution to the characterization of unlearning, we tested its inferential power, i.e. the capability of learning the original model. So far, to the best of our knowledge, the original unlearning algorithm was exclusively tested in an associative memory framework, where data are actually patterns, that are sub-dominant in number with respect to the entire set of possibile configurations. Our results show a good inferential performance of the algorithm. The explanation for this success relies on the choice of the initialization of the parameters. A Hebbian initialization of the network has two implications. First, the contribution to the gradient of the loss deriving from the data (i.e. the positive Hebbian contribution in the Boltzmann Machine learning in eq. 8) is much smaller than the contribution deriving from the cross-entropy of the model given the data (i.e. the negative unlearning contribution in the Boltzmann Machine learning). This condition holds until an optimal amount of iterations that can be estimated from the numerical simulations. Second, we showed that a Hebbian initialization of the couplings points the gradient of the parameters towards the correct direction, i.e. the one of the ground-truth of the inferential problem, at variance with other kinds of parameter initialization.
@@
We conclude that Boltzmann Machine learning can be performed in two steps: an initial overshooting along the direction of the data and then a gradual adjustment along the direction of the model with increments being updated at each step of the algorithm. The effectiveness of a two-step training for a Boltzmann Machine encourages a biological interpretation of this kind of training, in agreement with the hypothesis that synapses in the brain could be fine tuned through the alternation of daily online experience and offline sleep [36, 37]. Such a conjecture is inspiring members of the A.I. community in the development of training procedures for deep neural networks that might substitute the slow and non-biological back-propagation techniques [33, 38]. These observations, together with the importance of the Hebbian initialization of the network, appear to be consistent with previous results in literature concerning the unlearning algorithm and its implementation in associative memory modeling [10, 9, 11].

### v3

@@ V The Hebbian Unlearning limit
We will now study the specific limit of the unlearning regularization that leads to an algorithm which strongly resembles the HU routine. The inferential performance of the learning procedure is examined by measuring the distance between the inferred model and the original one at different training steps. Results show, for the first time, that HU can be employed as an inferential tool, due to the beneficial effect of the Hebbian initialization of the couplings. We advance an explanation for its performance that is supported by further numerical evidence. In the limit $a\rightarrow 0$ the updating rules for the parameters transform into Eqs.  (29) and (30). As mentioned before, Eq. (29) strongly resembles the traditional HU algorithm of Ref. [9]. By contrast with the original rule, the new regularization samples configurations at $\beta=1$ instead of stable fixed points of the neural dynamics, and makes use of a thermal average, rather than summing each contribution each time. We will keep dealing with small networks so that, given the initial conditions for the parameters, the loss function of the problem can be minimized exactly. As a numerical experiment, we train a BM with $N=18$ neurons to learn a SK model at $\beta_{d}=0.4$. The unlearning regularization is applied with $a=0$ and the usual initial conditions for the parameters, i.e. $\bm{J}^{(0)}=\bm{c}_{d}$ and $\bm{h}^{(0)}=\bm{0}$.
@@
VI A Real-World application
This section is dedicated to the testing of the performance of the Unlearning regularization on real data. We train a BM on a data-set available from Ref. [16], reporting the activity of $N=40$ retinal ganglion neurons reacting to visual stimuli. We firstly binarize each data vector by discretizing a time interval of $0.1s$ into bins $\Delta\tau=20ms$, as done in [16, 7]. As a consequence, each data-point is a sequence of $S_{i}$ units, representing the instance where the $i-th$ neuron has fired $(S_{i}=+1)$ or not fired $(S_{i}=-1)$ in given time bin. Each time bin is thus associated to a data-point of our binary data-set. We are provided with several repetitions of the experiment over the same population of neurons, so that we can train the model with $M=10^{5}$ data-points in total.
The model is trained as a BM and regularized over couplings and fields through the Unlearning method with different choices of the parameter $a$. The profile of the specific heat $C_{v}$ is measured as a function of $\beta$ and reported in Fig. 13. Lines with dots represent a machine trained until convergence of the algorithm. The curve obtained at $a=1$, corresponding to pure BM learning, appears consistent with the one reported in [7]. For each $a<1$ the peak of the specific heat shifts away from unit, and its value in $\beta=1$ is lowered by the regularization, suggesting a higher stability of the inferred network under a rescaling of couplings and fields. The small dots report the specific heat measured under the usual early-stopped prescription when couplings and fields are initialized as in equations (10), (11): even when $N$ becomes large there is an optimum amount of iterations where the Pearson coefficient between $\bm{c}_{d}$ and $\bm{c}_{1}$ is high and the slope $\omega$ reaches unity. As underlined by the figure, at this state of the network the specific heat is still significantly lower than the one obtained with no regularization, corroborating the importance of a Hebbian initialization for reaching the optimal level of accuracy and robustness of the model.
Figure 13: Specific heat of a model inferred from the activity of $N=40$ real neurons as a function of the inverse temperature $\beta$. The circles report the measure at convergence of the Unlearning regularization; the dots report measures when the algorithm is initialized in a Hebbian fashion and early-stopped when $\omega\simeq 1$. Error bars are neglected for clarity purposes. Choice of the parameters: $N=40$, $M=10^{5}$, $\lambda=10^{-3}$.
VII Conclusions
@@
The goal of this article was thus to show the effectiveness of unlearning as a form of regularization in Boltzmann Machines. The analysis has been carried out on both small networks, that allowed to compute all quantities without error by exact enumeration, and large networks, generally affected by statistical noise. For the most part of our work the ground-truth distribution was defined by one of two models, either the Curie-Weiss or the Sherrington-Kirkpatrick models, which are known to show a phase transition at a critical temperature. As a result, we could not only test the proximity of the inferred model to the original one, but also the susceptibility of the network under variations of the parameters, by comparing the trend of the specific heat of the ground truth model with respect to the inverse temperature, with the trend obtained after inference with regularization. In addition to this controlled framework, a real-world case, with data sampled from real neuronal activity, has been considered to confirm the results. Both heuristic and empirical arguments support the idea that models inferred via Boltzmann Machine learning tend to be pathologically critical [16, 7, 8], i.e. less effective in sampling new examples. On the other hand, the regularized model is much less critical than the one originally learned in [16] from the same data-set. Furthermore, we have displayed that even in the high-dimensional case one can characterize the performance of the algorithm in time by tracking two quantities: the Pearson coefficient and the slope of the empirical covariances versus the thermal correlations. These two observables are fast to compute and allow to establish an early stopping criterion that, combined with a Hebbian initialization of the parameters, reaches a greater generalization and robustness performance. We conclude that, given a high consistency between the model and the ground-truth distribution, the stability of the model under rescaling of the parameters is enhanced by the unlearning regularization. As an alternative strategy to the unlearning regularization presented here, we also evaluated the minimization of the KL distance between the models at $\beta=1$ and $\beta=a$. Even if the gradient descent equations look similar to the current case, they contain the contribution of four-point correlations among spins, resulting in a significant increase of the computational cost of the training.
@@
We also analyzed the particular limit of the regularization that gives as learning rules the ones expressed in equations (29) and (30). The resulting procedure strongly resembles the traditional Hebbian Unlearning routine, but with a thermal two-point correlation computed at $\beta=1$ replacing that computed on the zero-temperature fixed points of the dynamics. Previous work [13] supports the fact that the associative memory performance of such a thermal unlearning does not deviate much from the original rule. Unlearning is not new to be mentioned in the statistical inference realm [36, 37, 35], yet we are the first to consider the original rule proposed by Hopfield in [9]. As a new contribution to the characterization of unlearning, we tested its inferential power, i.e. the capability of learning the original model. So far, to the best of our knowledge, the original unlearning algorithm was exclusively tested in an associative memory framework, where data are actually patterns, that are sub-dominant in number with respect to the entire set of possibile configurations. Our results show a good inferential performance of the algorithm. The explanation for this success relies on the choice of the initialization of the parameters. A Hebbian initialization of the network has two implications. First, the contribution to the gradient of the loss deriving from the data (i.e. the positive Hebbian contribution in the Boltzmann Machine learning in eq. 8) is much smaller than the contribution deriving from the cross-entropy of the model given the data (i.e. the negative unlearning contribution in the Boltzmann Machine learning). This condition holds until an optimal amount of iterations that can be estimated from the numerical simulations. Second, we showed that a Hebbian initialization of the couplings points the gradient of the parameters towards the correct direction, i.e. the one of the ground-truth of the inferential problem, at variance with other kinds of parameter initialization. The beneficial effects of a Hebbian initialization for these type of machines has already emerged in the very recent literature [38, 12]. Yet, the topological properties of the Hebbian free energy landscape that make them such a good starting point for training have not received due attention.
@@
We conclude that Boltzmann Machine learning can be performed in two steps: an initial overshooting along the direction of the data and then a gradual adjustment along the direction of the model with increments being updated at each step of the algorithm. The effectiveness of a two-step training for a Boltzmann Machine encourages a biological interpretation of this kind of training, in agreement with the hypothesis that synapses in the brain could be fine tuned through the alternation of daily online experience and offline sleep [39, 40]. Such a conjecture is inspiring members of the A.I. community in the development of training procedures for deep neural networks that might substitute the slow and non-biological back-propagation techniques [35, 41]. These observations, together with the importance of the Hebbian initialization of the network, appear to be consistent with previous results in literature concerning the unlearning algorithm and its implementation in associative memory modeling [11, 10, 12, 38].